OpenStudy (anonymous):

A can of nuts sells for $4.50 and a box of microwaveable and a box of microwaveable popcorn sells for $3. You sell $252 in nuts and popcorn, and the boxes of popcorn sold is twice the amount of cans of nuts sold.How many boxes of popcorn did you sell?

OpenStudy (anonymous):

Equations: n=can of nuts; p=microwaveable popcorn 252=4.5n+3p (2n=p)-3 252=4.5n+3p + -6n=-3p 252-6n=4.5n +6n +6n (252=10.5n)/10.5 24=n 2(24)=p 48=p 252=4.5(24)+3(48) 252=108+144 252=252

OpenStudy (anonymous):

I usually get C's on all my Algebra tests cuz I'm working a grade above my grade level, but i got an A on this one :DD the only thing i actually understand!

OpenStudy (anonymous):

im confused on the second step you did------> (2n=p)-3

OpenStudy (anonymous):

@germanleopard ^

OpenStudy (anonymous):

that means \[(2n-p)\times3\]